OpenStudy (anonymous):

Write an equation of a line whose graph is parallel to the graph of y = 3x – 10.

OpenStudy (lgbasallote):

hint: the slope of that line is 3. so what do you think will be the equation of the parallel line?

OpenStudy (anonymous):

-3? @igbasallote

OpenStudy (lgbasallote):

nope. remember: parallel lines = same slopes

OpenStudy (anonymous):

Ah, sorry i guess i am not understanding

jimthompson5910 (jim_thompson5910):

y = mx+b y = 3x+b ... Plug in the slope m = 3 From here, you can make up any number for b and it will be your answer

OpenStudy (lgbasallote):

parallel lines have the same slopes perpendicular lines have negative reciprocal slopes

OpenStudy (lgbasallote):

what don't you understand?

OpenStudy (anonymous):

@jim_thompson5910, Igbasallote, so y=3x + 5? would be parallel then?

jimthompson5910 (jim_thompson5910):

yes, so are y = 3x+10 y = 3x+22 y = 3x - 16 ...etc etc

jimthompson5910 (jim_thompson5910):

as long as it has a slope of 3, it will be parallel

OpenStudy (anonymous):

Ok @jim_thompson5910, so my answer of y=3x + 5 would be correct for the question?

jimthompson5910 (jim_thompson5910):

yes it would be